Mike has $1.75 in dimes and nickels. If he has 5 more nickels than dimes, how many of each coin does he have?
? nickels
? dimes

Answer by algebrahouse.com(1659)

"Mike has $1.75 in dimes and nickels. If he has 5 more nickels than dimes, how many of each coin does he have?"

x = number of dimes
x + 5 = number of nickels {5 more nickels than dimes}

10x + 5(x + 5) = 175 {value of each x number of each = total value of each}
10x + 5x + 25 = 175 {used distributive property}
15x + 25 = 175 {combined like terms}
15x = 150 {subtracted 25 from both sides}
x = 10 {divided both sides by 15}
x + 5 = 15 {substituted 10, in for x, into x + 5}

10 dimes and 15 nickels
www.algebrahouse.com